# Biorthogonal measures, polymer partition functions, and random matrices

### v2

@@ 6 O’Connell-Yor polymer partition function and GUE with external source
Suppose that $|a_{i}|<\delta_{1}<1/2$ for all $i=1,\ldots,N$, and take $\delta_{2}\in(0,1)$ such that $\delta_{2}>2\delta_{1}$. Then, Borodin and Corwin [16] proved the identity
@@
    ${\rm d}\mu_{N}^{{\rm OY}}(\vec{x};\vec{a},\tau)=\frac{1}{Z_{N}}\det\left({\rm e}^{a_{m}x_{k}}\right)_{k,m=1}^{N}\det\left(\frac{1}{2\pi\mathrm{i}}\int_{\frac{1}{2}+\mathrm{i}\mathbb{R}}\frac{{\rm e}^{\tau\frac{t^{2}}{2}}{\rm e}^{-tx_{k}}}{(t-a_{m})\prod_{k=1}^{N}\Gamma(t-a_{k})}{\rm d}t\right)_{m,k=1}^{N}\prod_{k=1}^{N}{\rm d}x_{k}.$
@@
$\frac{1}{Z_{N}}{\rm e}^{-{\rm Tr}\,\left(\tau M^{2}/2-AM\right)}{\rm d}M,\quad{\rm d}M=\prod_{j=1}^{N}{\rm d}M_{jj}\ \prod_{1\leq j<k\leq N}{\rm d}\mathrm{Re}\,M_{jk}\,{\rm d}\mathrm{Im}\,M_{jk},$
@@
$\frac{1}{Z_{N}}\Delta(\vec{x})\ \det\left({\rm e}^{-\frac{\tau x_{j}^{2}}{2}+a_{k}x_{j}}\right)_{j,k=1}^{N}\prod_{j=1}^{N}{\rm d}x_{j},\qquad x_{1},\ldots,x_{N}\in\mathbb{R},$ (6.6)
@@ 7 Mixed polymer partition function
$(1,1)$$(\tau+N,N)$
$(s_{3},3)$$(s_{3},4)$$(s_{4},4)$$(s_{4},5)$$(s_{5},5)$$(s_{5},6)$$\alpha_{1}$$\alpha_{2}$$\alpha_{3}$$\alpha_{4}$$\alpha_{5}$$\alpha_{6}$
$a_{1}$$a_{2}$$a_{3}$$a_{4}$$a_{5}$$a_{6}$
@@
    ${\rm d}\mu_{N}^{{\rm Mixed}}(\vec{x};\vec{\alpha},\vec{a},\tau)=\frac{1}{Z_{N}}\det\left(x_{k}^{m-1}\right)_{k,m=1}^{N}\det\left(\frac{1}{2\pi i}\int_{\frac{1}{2}+\mathrm{i}\mathbb{R}}\frac{{\rm e}^{\tau\frac{t^{2}}{2}}\Gamma(\alpha-t)^{n}{{\rm e}}^{-(t-a)x_{k}}}{(t-a)^{N-m+1}\Gamma(t-a)^{N}}\mathrm{d}t\right)_{m,k=1}^{N}\prod_{k=1}^{N}{\rm d}x_{k}.$
@@
Consider an $N\times N$ matrix $M$ from the LUE with external source, with distribution (5.11) and $\nu=0$, and consider an $N\times N$ GUE matrix $H$, with distribution (6.6) with $A=0$. Set $Q=M+\sqrt{\tau N}H$. The eigenvalue distribution of $Q$ can be computed using [27, Theorem 2.3 (b)], or more directly using the re-scaled variant of this identity given in [25, Equation (1.6)]. This result gives an integral expression for the correlation kernel of the eigenvalues of $Q$ in terms of the eigenvalue correlation of $M$, valid for any random matrix $M$ whose eigenvalue distribution is a polynomial ensemble, and thus in particular for our LUE with external source random matrix $M$. It states that the eigenvalue correlation kernel $K_{N}(x,x^{\prime})$ of $Q$ is given by
$K_{N}(x,x^{\prime})=\frac{1}{2\pi\mathrm{i}\tau}\int_{\mathrm{i}\mathbb{R}}{\rm d}s\int_{\mathbb{R}}{\rm d}t\;k_{N}(s,t){\rm e}^{\frac{1}{2\tau}\left((x-s)^{2}-(x^{\prime}-t)^{2}\right)},$ (7.11)
where
$k_{N}(s,t)=\sum_{k=1}^{N}p_{k}(s)q_{k}(t),$
with $p_{k}$ a polynomial of degree $k-1$, $q_{1},\ldots,q_{N}$ in the linear span of the functions $g_{1},\ldots,g_{N}$ from (1.1), and they are biorthogonal:
@@
In our situation, we can take $k_{N}(s,t)$ of the form (1.8) with $W_{N}=W_{N}^{\rm LUE+}$ as in (5.14) with $\nu=0$. Then $K_{N}(x,x^{\prime})$ is given by
$K_{N}(x,x^{\prime})$ $=\frac{1}{2\pi\mathrm{i}\tau}\int_{\mathrm{i}\mathbb{R}}{\rm d}s\int_{\mathbb{R}}{\rm d}t\;L_{N}^{{\rm LUE+}}(s,t{;\vec{a},0}){\rm e}^{\frac{1}{2\tau}\left((x-s)^{2}-(x^{\prime}-t)^{2}\right)}$
$=\frac{1}{(2\pi\mathrm{i})^{3}\tau}\int_{\mathrm{i}\mathbb{R}}{\rm d}s\int_{\mathbb{R}}{\rm d}t\int_{\Sigma_{N}}{\rm d}u\int_{\ell_{N}}{\rm d}v\frac{W_{N}^{{\rm LUE+}}(v{;\vec{a},0})}{W_{N}^{{\rm LUE+}}(u{;\vec{a},0})}\frac{{\rm e}^{-vs+ut}}{v-u}{\rm e}^{\frac{1}{2\tau}\left((x-s)^{2}-(x^{\prime}-t)^{2}\right)}.$
Changing the order of integration, we obtain after a straightforward computation that
$K_{N}(x,x^{\prime})$ $=\frac{1}{(2\pi\mathrm{i})^{3}\tau}\int_{\Sigma_{N}}{\rm d}u\int_{\ell_{N}}{\rm d}v\frac{W_{N}^{{\rm LUE+}}(v{;\vec{a},0})}{W_{N}^{{\rm LUE+}}(u{;\vec{a},0})}\frac{{\rm e}^{\frac{\tau}{2}v^{2}-\frac{\tau}{2}u^{2}+ux^{\prime}-vx}}{v-u}$
×$\qquad\qquad\qquad\times\ \left(\int_{\mathrm{i}\mathbb{R}}{\rm d}s\;{\rm e}^{\frac{1}{2\tau}\left(s-(x+\tau v)\right)^{2}}\right)\left(\int_{\mathbb{R}}{\rm d}t\;{\rm e}^{-\frac{1}{2\tau}\left(t-(x^{\prime}+\tau v)\right)^{2}}\right)$
$=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma_{N}}{\rm d}u\int_{\ell_{N}}{\rm d}v\frac{W_{N}^{{\rm LUE+}}(v{;\vec{a},0})}{W_{N}^{{\rm LUE+}}(u{;\vec{a},0})}\frac{{\rm e}^{\frac{\tau}{2}v^{2}-\frac{\tau}{2}u^{2}+ux^{\prime}-vx}}{v-u}.$
This is again of the form (1.8), but now with transformed function $W$ given by where $W_{N}^{\rm GLUE+}$ is defined by
$W_{N}^{\rm GLUE+}(v;\vec{a},\tau):=\frac{{\rm e}^{\tau z^{2}/2}\prod_{k=1}^{N}(z-a_{k})}{(1-z)^{N}}.$ (7.12)
@@
Consider the mixed polymer partition function with re-scaled $\tau=\tau/T^{2}$, re-scaled $a_{m}$-parameters $Ta_{1},\ldots,Ta_{N}$, and $\alpha_{1}=\cdots=\alpha_{N}={T}$, in the limit where $T\to 0$. Then, the kernel of the associated biorthogonal measure $L_{N}$ is given by
$L_{N}^{\rm Mixed}\left(x,x^{\prime};\left[T\right]_{N},T\vec{a},\tau/T^{2}\right)=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma_{N}}{\rm d}u\int_{\ell_{N}}{\rm d}v\frac{W_{N}^{\rm Mixed}(v;\left[T\right]_{N},T\vec{a},\tau/T^{2})}{W_{N}^{\rm Mixed}(u;\left[T\right]_{N},T{\vec{a}},\tau/T^{2})}\frac{{\rm e}^{-vx+ux^{\prime}}}{v-u},$
@@
$\frac{1}{T}L_{N}^{\rm Mixed}(\frac{x}{T},\frac{x^{\prime}}{T};\left[T\right]_{N},T\vec{a},\tau/T^{2})=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\widetilde{\Sigma}_{N}}{\rm d}s\int_{\widetilde{\ell}_{N}}{\rm d}t\frac{\widetilde{W}_{N}^{\rm Mixed}(t;[1]_{N},{\vec{a}},\tau)}{\widetilde{W}_{N}^{\rm Mixed}(s;[1]_{N},{\vec{a}},\tau,T)}\frac{{\rm e}^{-tx+sx^{\prime}}}{t-s},$
@@
$\widetilde{W}_{N}^{\rm Mixed}(z;[1]_{N},\vec{a},\tau)=\frac{{\rm e}^{\tau z^{2}/2}\Gamma((1-z)T)^{N}}{\prod_{k=1}^{N}\Gamma((z-a_{k})T)}=\frac{\Gamma(1+(1-z)T)^{N}\prod_{k=1}^{N}(z-a_{k})}{(1-z)^{N}\prod_{k=1}^{N}\Gamma(1+(z-a_{k})T)}{\rm e}^{\tau z^{2}/2},$
@@
$\lim_{T\to 0}\frac{\widetilde{W}_{N}^{\rm Mixed}(v;[1]_{N},{\vec{a}},\tau)}{\widetilde{W}_{N}^{\rm Mixed}(u;[1]_{N},{\vec{a}},\tau)}=\frac{W_{N}^{\rm GLUE+}(v;{\vec{a}},\tau)}{W_{N}^{\rm GLUE+}(u;{\vec{a}},\tau)}.$
@@
$\lim_{T\to 0}\frac{1}{T}L_{N}^{{\rm Mixed}}\left(\frac{x}{T},\frac{x^{\prime}}{T};\left[T\right]_{N},T{\vec{a}},\tau/T^{2}\right)=L_{N}^{\rm GLUE+}\left(x,x^{\prime};\vec{a},\tau\right),\qquad x,x^{\prime}\in\mathbb{R},$ (7.13)
and there exists $C_{N}(\vec{a},\tau)$ such that for $\epsilon>0$ sufficiently small and $T>0$ sufficiently small,
$\left|\frac{1}{T}L_{N}^{{\rm Mixed}}\left(\frac{x}{T},\frac{x^{\prime}}{T};\left[T\right]_{N},T{\vec{a}},\tau/T^{2}\right)\right|\leq C_{N}(\vec{a},\tau)h(x)\tilde{h}(x^{\prime}),$ (7.14)
with $-~h,\tilde{h}$ as in (5.18).
@@
$\mathbb{E}\left[{\rm e}^{-{\rm e}^{t/T}Z_{N,N}^{{\rm Mixed}}\left(\left[T\right]_{N},T{\vec{a}},\tau/T^{2}\right)}\right]\longrightarrow\mathbb{P}_{\mathrm{GLUE+}}\left(\max\{x_{1},\ldots,x_{N}\}\leq-t;\vec{a},\tau\right),$
@@
GUE $z^{N}{\rm e}^{\tau z^{2}/2}$ $x^{m-1}$ $x^{m-1}{\rm e}^{-\tau x^{2}/2}$
GUE + external source $\prod_{m=1}^{N}(z-a_{m})\ {\rm e}^{\tau z^{2}/2}$ ${x^{m-1}}$ ${{{\rm e}}^{-(\tau x^{2}/2-a_{m}x)}}$
@@
LUE + external source $\frac{\prod_{m=1}^{N}(z-a_{m})}{(z-1)^{N+\nu}}$ $x^{m-1}$ $x^{\nu}{{\rm e}}^{-(1-a_{m})x}$
(LUE + external  source) + GUE $\frac{{\rm e}^{\tau z^{2}/2}\prod_{k=1}^{N}(z-a_{k})}{(1-z)^{N}}$ $x^{m-1}$ $\frac{1}{2\pi\mathrm{i}}\int_{\mathrm{i}\mathbb{R}}\frac{{\rm e}^{\tau\frac{t^{2}}{2}}\prod_{j=1}^{N}(t-a_{j})}{(t-a_{m})(t-1)^{N}}{{\rm e}}^{-tx}\mathrm{d}t$
@@
Table 1: List of polymer and random matrix models associated to biorthogonal measures with symbol $F(z)$, and a possible choice, not necessarily biorthogonal, for the corresponding functions $f_{m}(x)$ and $g_{m}(x)$ in (1.1), for $m=1,\ldots,N$. The choices for $g_{m}$ in the GUE with external source and LUE with external source are only valid if $a_{1},\ldots,a_{N}$ are distinct. The last three models correspond to the case $a_{m}=m-1$, and the eigenvalues or squared singular values in these models correspond to the exponential variable $y={\rm e}^{x}$.

### v3

@@ 6 O’Connell-Yor polymer partition function and GUE with external source
Suppose that $|a_{i}|<\delta_{1}<1/2$ for all $i=1,\ldots,N$, and take $\delta_{2}\in(0,1)$ such that $\delta_{2}>2\delta_{1}$. Then, Borodin and Corwin [16, Theorem 5.2.11] proved the identity
@@
    ${\rm d}\mu_{N}^{{\rm OY}}(\vec{x};\vec{a},\tau)=\frac{1}{Z_{N}}\det\left({\rm e}^{a_{m}x_{k}}\right)_{k,m=1}^{N}\det\left(\frac{1}{2\pi\mathrm{i}}\int_{\frac{1}{2}+\mathrm{i}\mathbb{R}}\frac{{\rm e}^{\tau\frac{t^{2}}{2}}{\rm e}^{-tx_{k}}}{(t-a_{m}){\prod_{j=1}^{N}\Gamma(t-a_{j})}}{\rm d}t\right)_{m,k=1}^{N}\prod_{k=1}^{N}{\rm d}x_{k}.$
@@
Remark 6.2.
Using the integral representations of $\{\psi_{i}\}_{i=1}^{N}$ and $\{\phi_{i}\}_{i=1}^{N}$ with $W_{N}$ given by (6.3), it is easy to prove that
$\psi_{i}({\rm e}^{\bar{x}})=\overline{\psi_{i}({\rm e}^{x})},\quad\phi_{i}({\rm e}^{\bar{x}})=\overline{\phi_{i}({\rm e}^{x})},\quad i=1,\ldots,N.$
Consequently, $d\mu_{N}^{{\rm OY}}$ is a signed real-valued measure.
@@
$\frac{1}{Z_{N}}{\rm e}^{-{\rm Tr}\,\left(\frac{M^{2}}{2\tau}-AM\right)}{\rm d}M,\quad{\rm d}M=\prod_{j=1}^{N}{\rm d}M_{jj}\ \prod_{1\leq j<k\leq N}{\rm d}\mathrm{Re}\,M_{jk}\,{\rm d}\mathrm{Im}\,M_{jk},$
@@
$\frac{1}{Z_{N}}\Delta(\vec{x})\ \det\left({\rm e}^{-\frac{x_{j}^{2}}{2\tau}+a_{k}x_{j}}\right)_{j,k=1}^{N}\prod_{j=1}^{N}{\rm d}x_{j},\qquad x_{1},\ldots,x_{N}\in\mathbb{R},$ (6.6)
@@ 7 Mixed polymer partition function
$(1,1)$$(\tau+N,N)$$(s_{3},3)$$(s_{3},4)$$(s_{4},4)$$(s_{4},5)$$(s_{5},5)$$(s_{5},6)$$\alpha_{1}$$\alpha_{2}$$\alpha_{3}$$\alpha_{4}$$\alpha_{5}$$\alpha_{6}$$a_{1}$$a_{2}$$a_{3}$$a_{4}$$a_{5}$$a_{6}$
@@
    ${\rm d}\mu_{N}^{{\rm Mixed}}(\vec{x};\vec{\alpha},\vec{a},\tau)=\frac{1}{Z_{N}}\det\left(x_{k}^{m-1}\right)_{k,m=1}^{N}\det\left(\frac{1}{2\pi i}\int_{\frac{1}{2}+\mathrm{i}\mathbb{R}}\frac{{\rm e}^{\tau\frac{t^{2}}{2}}{\prod_{j=1}^{n}\Gamma(\alpha_{j}-t)}{{\rm e}}^{-(t-a)x_{k}}}{(t-a)^{N-m+1}\Gamma(t-a)^{N}}\mathrm{d}t\right)_{m,k=1}^{N}\prod_{k=1}^{N}{\rm d}x_{k}.$
Remark 7.4.
Using the integral representations of $\{\psi_{i}\}_{i=1}^{N}$ and $\{\phi_{i}\}_{i=1}^{N}$ with $W_{N}$ given by (7.8), it is easy to prove that
$\psi_{i}({\rm e}^{\bar{x}})=\overline{\psi_{i}({\rm e}^{x})},\quad\phi_{i}({\rm e}^{\bar{x}})=\overline{\phi_{i}({\rm e}^{x})},\quad i=1,\ldots,N.$
Consequently, $d\mu_{N}^{{\rm Mixed}}$ is a signed real-valued measure.
@@
Consider an $N\times N$ matrix $M$ from the LUE with external source, with distribution (5.11) and $\nu=0$, and consider an $N\times N$ GUE matrix $H$, with distribution (6.6) with $A=0$. Set $Q=M+\sqrt{\tau N}H$. The eigenvalue distribution of $Q$ can be computed using [27, Theorem 2.3 (b)], or more directly using the re-scaled variant of this identity given in [25, Equation (1.6)]. This result gives an integral expression for the correlation kernel of the eigenvalues of $Q$ in terms of the eigenvalue correlation of $M$, valid for any random matrix $M$ whose eigenvalue distribution is a polynomial ensemble, and thus in particular for our LUE with external source random matrix $M$. It states that an eigenvalue correlation kernel $K_{N}(x,x^{\prime})$ of $Q$ is given by
$K_{N}(x,x^{\prime})=\frac{1}{2\pi\mathrm{i}\tau}\int_{\mathrm{i}\mathbb{R}}{\rm d}s\int_{\mathbb{R}}{\rm d}t\;k_{N}(s,t){\rm e}^{\frac{1}{2\tau}\left((x^{\prime}-s)^{2}-(x-t)^{2}\right)},$ (7.11)
where $k_{N}(s,t)$ is a correlation kernel of the eigenvalue distribution of $M$ which is polynomial in $s$: more precisely, it is of the form
$k_{N}(s,t)=\sum_{k=1}^{N}p_{k}(s)q_{k}(t),$ (7.12)
where $p_{1},\ldots,p_{N}$ are polynomials of degree $N-1$, and we have the biorthogonality relations
@@
In our situation, we can take
$k_{N}(s,t)=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma}{\rm d}u\int_{\ell}{\rm d}v\left(\frac{v}{u}\right)^{N}\prod_{j=1}^{N}\frac{u-1+b_{j}}{v-1+b_{j}}\frac{{\rm e}^{us-vt}}{v-u},\qquad s,t\in\mathbb{R}$ (7.13)
where $\Sigma$ encircles 00, and $\ell=c+\mathrm{i}\mathbb{R}$ is a vertical line at the right of $\Sigma$ and at the left of $1-b_{1},\ldots,1-b_{N}$. The $v$-integral is defined as a Cauchy principal value integral, $\int_{\ell}=\lim_{R\to\infty}\int_{c-\mathrm{i}R}^{c+\mathrm{i}R}{\rm d}v$. Indeed, for $t<0$, we can compute this principal value integral by closing the integration contour $[c-\mathrm{i}R,c+\mathrm{i}R]$ with a large semi-circle to the left, and it follows then from a residue computation that $k_{N}(s,t)=0$. Similarly, for $t>0$, we can close the integration contour with a large semi-circle on the right, and then we obtain that
$k_{N}(s,t)={\rm e}^{s-t}L_{N}(s,t),$
with $L_{N}=L_{N}^{\rm LUE+}$ given by (5.12) with $\nu=0$. It is also straightforward to verify using the residue theorem that $k_{N}(s,t)$ is polynomial in $s$, and more precisely of the required form (7.12). As for the biorthogonality relations, we refer to [10]. We then apply (7.11) and find that $K_{N}(x,x^{\prime})$ is given by
$K_{N}(x,x^{\prime})$ $=\frac{1}{2\pi\mathrm{i}\tau}\int_{\mathrm{i}\mathbb{R}}{\rm d}s\int_{\mathbb{R}}{\rm d}t\;k_{N}(s,t){\rm e}^{\frac{1}{2\tau}\left((x^{\prime}-s)^{2}-(x-t)^{2}\right)}$
$=\frac{1}{(2\pi\mathrm{i})^{3}\tau}\int_{\mathrm{i}\mathbb{R}}{\rm d}s\int_{\mathbb{R}}{\rm d}t\int_{\Sigma}{\rm d}u\int_{\ell}{\rm d}v\left(\frac{v}{u}\right)^{N}\prod_{j=1}^{N}\frac{u-1+b_{j}}{v-1+b_{j}}\frac{{\rm e}^{us-vt}}{v-u}{\rm e}^{\frac{1}{2\tau}\left((x^{\prime}-s)^{2}-(x-t)^{2}\right)}.$
Changing the order of integration and writing $\tilde{W}_{N}(z)=\frac{z^{N}}{\prod_{j=1}^{N}(z-1+b_{j})}$, we obtain after a straightforward computation that
$K_{N}(x,x^{\prime})$ $=\frac{1}{(2\pi\mathrm{i})^{3}\tau}\int_{\Sigma}{\rm d}u\int_{\ell}{\rm d}v\frac{\tilde{W}_{N}(v)}{\tilde{W}_{N}(u)}\frac{{\rm e}^{\frac{\tau}{2}v^{2}-\frac{\tau}{2}u^{2}+ux^{\prime}-vx}}{v-u}$
×$\qquad\qquad\qquad\times\ \left(\int_{\mathrm{i}\mathbb{R}}{\rm d}s\;{\rm e}^{\frac{1}{2\tau}\left(x^{\prime}-s-\tau u\right)^{2}}\right)\left(\int_{\mathbb{R}}{\rm d}t\;{\rm e}^{-\frac{1}{2\tau}\left(x-t-\tau v\right)^{2}}\right)$
$=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma}{\rm d}u\int_{\ell}{\rm d}v\frac{\tilde{W}_{N}(v){\rm e}^{\frac{\tau}{2}v^{2}}}{\tilde{W}_{N}(u){\rm e}^{\frac{\tau}{2}u^{2}}}\frac{{\rm e}^{x^{\prime}u-xv}}{v-u}.$
We conclude that $K_{N}(x,x^{\prime})$ is again a kernel of the form (1.8), but now with transformed function $W$ given by
$W_{N}^{\rm GLUE+}(z;\vec{b},\tau):=\frac{z^{N}{\rm e}^{\tau z^{2}/2}}{\prod_{k=1}^{N}(z-1+b_{k})}.$ (7.14)
@@
Consider the mixed polymer partition function with re-scaled $\tau=\tau/T^{2}$, $a_{1}=\cdots=a_{N}=0$, and with re-scaled $\alpha$-parameters $\alpha_{1}=T-Tb_{1},\ldots,\alpha_{N}=1-Tb_{N}$, in the limit where $T\to 0$. Then, the kernel of the associated biorthogonal measure $L_{N}$ is given by
${L_{N}^{\rm Mixed}\left(x,x^{\prime};\left[T\right]_{N}-T\vec{b},[0]_{N},\tau/T^{2}\right)}=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma_{N}}{\rm d}u\int_{\ell_{N}}{\rm d}v\frac{W_{N}^{\rm Mixed}(v;[T]_{N}-T\vec{b},[0]_{N},\tau/T^{2})}{W_{N}^{\rm Mixed}(u;[T]_{N}-T\vec{b},[0]_{N},\tau/T^{2})}\frac{{\rm e}^{-vx+ux^{\prime}}}{v-u},$
@@
$\frac{1}{T}L_{N}^{\rm Mixed}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{N}-T\vec{b},[0]_{N},\tau/T^{2}\right)=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\widetilde{\Sigma}_{N}}{\rm d}s\int_{\widetilde{\ell}_{N}}{\rm d}t\frac{\widetilde{W}_{N}^{\rm Mixed}(t;[1]_{N}-\vec{b},[0]_{N},\tau,T)}{\widetilde{W}_{N}^{\rm Mixed}(s;[1]_{N}-\vec{b},[0]_{N},\tau,T)}\frac{{\rm e}^{-tx+sx^{\prime}}}{t-s},$
@@
$\widetilde{W}_{N}^{\rm Mixed}(z;[1]_{N}-\vec{b},[0]_{N},\tau,T)=\frac{{\rm e}^{\tau z^{2}/2}\prod_{k=1}^{N}\Gamma((1-b_{k}-z)T)}{\Gamma(zT)^{N}}=\frac{z^{N}{\rm e}^{\tau z^{2}/2}}{\prod_{k=1}^{N}(1-b_{k}-z)}\frac{\prod_{k=1}^{N}\Gamma(1+(1-b_{k}-z)T)}{\Gamma(1+zT)^{N}},$
@@
$\lim_{T\to 0}\frac{\widetilde{W}_{N}^{\rm Mixed}(v;[T]_{N}-T\vec{b},[0]_{N},\tau,T)}{\widetilde{W}_{N}^{\rm Mixed}(u;[T]_{N}-T\vec{b},[0]_{N},\tau,T)}=\frac{W_{N}^{\rm GLUE+}(v;\vec{b},\tau)}{W_{N}^{\rm GLUE+}(u;\vec{b},\tau)}.$
@@
$\lim_{T\to 0}\frac{1}{T}L_{N}^{{\rm Mixed}}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{N}-T\vec{b},[0]_{N},\tau/T^{2}\right)=L_{N}^{\rm GLUE+}\left(x,x^{\prime};\vec{b},\tau\right),\qquad x,x^{\prime}\in\mathbb{R},$ (7.15)
and there exists $C_{N}(\vec{b},\tau)$ such that for $\epsilon>0$ sufficiently small and $T>0$ sufficiently small,
$\left|\frac{1}{T}L_{N}^{{\rm Mixed}}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{N}-T\vec{b},[0]_{N},\tau/T^{2}\right)\right|\leq C_{N}(\vec{b},\tau)h(x)\tilde{h}(x^{\prime}),$ (7.16)
with $h,\tilde{h}$ as in (5.18).
@@
$\mathbb{E}\left[{\rm e}^{-{\rm e}^{t/T}Z_{N,N}^{{\rm Mixed}}\left([T]_{N}-T\vec{b},[0]_{N},\tau/T^{2}\right)}\right]\longrightarrow\mathbb{P}_{\mathrm{GLUE+}}\left(\max\{x_{1},\ldots,x_{N}\}\leq-t;\vec{b},\tau\right),$
@@
GUE $z^{N}{\rm e}^{\tau z^{2}/2}$ $x^{m-1}$ $x^{m-1}{\rm e}^{-\frac{x^{2}}{2\tau}}$
GUE + external source $\prod_{m=1}^{N}(z-a_{m})\ {\rm e}^{\tau z^{2}/2}$ ${x^{m-1}}$ ${{{\rm e}}^{-(\frac{x^{2}}{2\tau}-a_{m}x)}}$
@@
LUE + external source $\frac{\prod_{m=1}^{N}(z-b_{m})}{(z-1)^{N+\nu}}$ $x^{m-1}$ $x^{\nu}{{\rm e}}^{-(1-b_{m})x}$
(LUE + external  source) + GUE $\frac{{\rm e}^{\tau z^{2}/2}z^{N}}{\prod_{k=1}^{N}(z-1+b_{k})}$ $x^{m-1}$ $\frac{1}{2\pi\mathrm{i}}\int_{\mathrm{i}\mathbb{R}}\frac{{\rm e}^{\tau\frac{t^{2}}{2}}t^{m-1}}{\prod_{j=1}^{N}(t-1+b_{j})}{{\rm e}}^{-tx}\mathrm{d}t$
@@
Table 1: List of polymer and random matrix models associated to biorthogonal measures with symbol $F(z)$, and a possible choice, not necessarily biorthogonal, for the corresponding functions $f_{m}(x)$ and $g_{m}(x)$ in (1.1), for $m=1,\ldots,N$. The choices for $g_{m}$ in the GUE with external source and LUE with external source are only valid if $a_{1},\ldots,a_{N}$ respectively $b_{1},\ldots,b_{N}$ are distinct. The last three models correspond to the case $a_{m}=m-1$, and the eigenvalues or squared singular values in these models correspond to the exponential variable $y={\rm e}^{x}$.